# F-95000 Cergy-Pontoise, France

### v1

@@ Appendix C Proof of Proposition 10
  • For $aggr=sum$, $min\_ans$, respectively $max\_ans$, is the minimal, respectively maximal, sum that can be obtained by adding one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$ and then possibly one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed in lines 17 and 30, respectively in lines 18 and 32 of Algorithm 2. Notice that if all $M_{i}$-values are positive, knowing that adding 00 is neutral, then the test line 29 always fails and thus, $proc\_min$ is the sum of all minimal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$. Similarly, in this case, the test line 31 always succeeds, except for 00, and thus, $proc\_max$ is the sum of all maximal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma(\mathcal{AQ})$.
@@
To complete the proof that the returned values $min\_ans$ and $max\_ans$ are as stated in the proposition, we have to show that ${\sf Rep}(T)$ contains repairs $R_{min}$ and $R_{max}$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{min}]})=min\_ans$ and ${\sf Ans}(\mathcal{AQ}^{[R_{max}]})=max\_ans$. These results are consequences of Proposition 9 by considering $R_{min}=R_{2}$ and $R_{max}=R_{1}$ if $aggr=min$, and $R_{min}=R_{1}$ and $R_{max}=R_{2}$ if $aggr$ is $max$ or $count$ or if $M_{i}$-values are positive and $aggr=sum$. We also mention that if $aggr=sum$ with no restriction on the $M_{i}$-values, $min\_ans$, respectively $max\_ans$, is the least, respectively the largest, sum that can be obtained using all relevant sets $\sigma(M_{i})$. We thus have $min\_ans\leq glb$ and $max\_ans\geq lub$, which concludes this part of the proof.

### v2

@@ Appendix G Proof of Proposition 11
  • For $aggr=sum$, min_ansmin\_ans, respectively max_ansmax\_ans, is the minimal, respectively maximal, sum that can be obtained by adding one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$ and then possibly one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed in lines 18 and 30, respectively in lines 18 and 32 of Algorithm 3. Notice that if all $M_{i}$-values are positive, knowing that adding 0 is neutral, then the test line 29 always fails and thus, proc_minproc\_min is the sum of all minimal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$. Similarly, in this case, the test line 31 always succeeds, except for 0, and thus, proc_maxproc\_max is the sum of all maximal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma(\mathcal{AQ})$.
@@
To complete the proof that the returned values min_ansmin\_ans and max_ansmax\_ans are as stated in the proposition, we have to show that ${\sf Rep}(T)$ contains repairs $R_{min}$ and $R_{max}$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{min}]})=min\_ans$ and ${\sf Ans}(\mathcal{AQ}^{[R_{max}]})=max\_ans$. These results are consequences of Proposition 10 by considering $R_{min}=R_{2}$ and $R_{max}=R_{1}$ if $aggr=min$, and $R_{min}=R_{1}$ and $R_{max}=R_{2}$ if aggraggr is maxmax or countcount or if $M_{i}$-values are positive and $aggr=sum$. We also mention that if $aggr=sum$ with no restriction on the $M_{i}$-values, min_ansmin\_ans, respectively max_ansmax\_ans, is the least, respectively the largest, sum that can be obtained using all relevant sets $\sigma(M_{i})$. We thus have $min\_ans\leq glb$ and $max\_ans\geq lub$, which concludes this part of the proof.